OpenStudy (anonymous):

If y varies jointly as x and the cube root of z, and y = 120 when x = 3 and z = 8, find y when x = 4 and z = 27

OpenStudy (anonymous):

y = kxz^(1/3) 120 = k(3)(8)^(1/3) 120 = k(3)(2) k = 20 y = 20xz^(1/3) y = 20(4)27^(1/3) y = 240

OpenStudy (anonymous):

All good?

OpenStudy (anonymous):

The trick is the very first equation and whether inverse or not.

OpenStudy (anonymous):

Here, we had a direct proportion whereas in the previous problem, there was an inverse proportionality, hence the denominator in the first answer of the 2 questions.

OpenStudy (anonymous):

If you go over these 2 questions as a model, you'll be able to do all future similar questions with no sweat.

OpenStudy (anonymous):

And of course "k" is essential.